# Bistability and chaos in the discrete two-gene Andrecut-Kauffman model

### v1

@@ 4. An absorbing set for the system
Following the ideas from [PG24], we shall prove that if $x_{t}\geq b_{1}$ then $x_{t+1}\leq x_{t}-c_{1}$, that is, after each iteration the $x$ coordinate of the point decreases at least by the constant amount $c_{1}$, as long as it is beyond $b_{1}$. This statement implies that after a finite number $N_{0}$ of iterations, we must reach the condition $x_{t+N_{0}}\leq b_{1}$. Moreover, we shall also show that if $x_{t}\leq b_{1}$ then also $x_{t+1}\leq b_{1}$.
@@
(3) $x_{t+1}=\alpha_{1}/(1+\text{a non-negative term})+\beta x_{t}\leq\alpha_{1}+\beta x_{t}\\ \leq(1-\beta)x_{t}-c_{1}+\beta x_{t}=x_{t}-c_{1}.$
Let us consider the second statement now, and take $x_{t}\leq b_{1}=(\alpha_{1}+c_{1})/(1-\beta)$. Then we calculate:
(4) $x_{t+1}\leq\alpha_{1}+\beta x_{t}\leq\alpha_{1}+\beta(\alpha_{1}+c_{1})/(1-\beta)\\ =\alpha_{1}(1-\beta)/(1-\beta)+\beta(\alpha_{1}+c_{1})/(1-\beta)\\ =(\alpha_{1}-\beta\alpha_{1}+\beta\alpha_{1}+\beta c_{1})/(1-\beta)\\ =(\alpha_{1}+\beta c_{1})/(1-\beta)\leq(\alpha_{1}+c_{1})/(1-\beta)=b_{1}.$
@@
For example, it follows from Proposition 2 that if $\alpha_{i}\in[0,100]$ and $\beta\leq 0.5$ then all the recurrent dynamics of the system we consider is contained in $[0,200]^{2}$.
@@ 5. Periodic and chaotic behavior of the model
Both models (1) and (2) admit periodic solutions and chaotic dynamics. However, allowing $\alpha_{1}\neq\alpha_{2}$ may switch the system to quantitatively different dynamics by changing $\alpha_{2}$ alone while leaving $\alpha_{1}$ unchanged, as we illustrate below.
The plots in Figure 3 show time evolution of the variables $x$ and $y$ in the system (2) for two specific trajectories, both starting at $x_{0}=y_{0}=0.5$: one for the system with $\alpha_{1}=\alpha_{2}=20$, the other one with the same value of the first reaction rate $\alpha_{1}=20$, but with $\alpha_{2}=60$. In this example, the same values of the other parameters are used in both cases: $\varepsilon=0.8$, $\beta=0.1$, $n=4$. For $\alpha_{1}=\alpha_{2}=20$ one can immediately notice the chaotic nature of the system. However, the trajectory for $\alpha_{2}=60$ appears to approach a stable periodic orbit.
The opposite situation is also present in the model. Namely, it is possible to find a case in which for $\alpha_{1}=\alpha_{2}$ we observe periodic solutions, while changing the value of $\alpha_{2}$ yields chaotic behavior. An example of a such a case occurs for the parameters $\alpha_{1}=3$, $\alpha_{2}=3$, $\beta=1$, $\varepsilon=0.8$, and $n=4$. Periodic trajectories switch to chaotic solutions when we change the value of $\alpha_{2}$ to $28$.
These plots highlight the role of $\alpha_{1}$ and $\alpha_{2}$ in controlling the system’s stability, and illustrate how the gene expression model can shift between oscillatory and chaotic regimes, depending solely on the choice of these parameters.
@@ 6. Analysis of bifurcation diagrams
A bifurcation diagram is a visual representation of the evolution of the global attractor (or attractors in general) of a dynamical system as a selected one-dimensional parameter varies. The horizontal axis of the diagram typically represents the parameter being varied, while the vertical axis corresponds to the one-dimensional phase space of the system, or a projection onto a selected variable if the actual dimension of the phase space is higher than one. Several features of the global attractor can be easily spotted on a bifurcation diagram, such as the size and the complexity of the attractor, as well as bifurcations it undergoes when the parameter is varied.
The purpose of the experiment described below is to investigate the dynamics by means of bifurcation diagrams with respect to one of the reaction rates $\alpha_{1}$ while keeping the other rate $\alpha_{2}$ fixed in the system (2) and compare it to the the dynamics of the system (1) in which both reaction rates are the same. For that end, we fix $\varepsilon=0.7$, $\beta=0.2$, $n=4$, $\alpha_{2}=200$, and we compute bifurcation diagrams for $\alpha_{1}\in[0,500]$ for the system (2), and we do the same for the system (1) with $\alpha\in[0,500]$.
The results of this experiment are shown in Figure 4. In the vertical axes of the bifurcation diagrams, we show the projections of the observed dynamics to the $x$ and $y$ variables, respectively. In the horizontal axes, we indicate the parameters of the systems that are varied: the common reaction rate $\alpha$ of the system (1) in Figures 4(a) and 4(c), and the reaction rate $\alpha_{1}$ of the system (2) in Figures 4(b) and 4(b).
It is interesting to see that there is a sudden transition in the system (2) that occurs in the bifurcation diagrams 4(b) and 4(d) as the parameter $\alpha_{1}$ crosses the value $\alpha_{2}=200$. Indeed, this is kind of a boundary between two different behaviors of the model. Additionally, we notice considerable difference between the shape of the diagram 4(b) and 4(d). The first one resembles a series of period doubling bifurcations, while the second one looks like showing period-halving bifurcations.
On the other hand, the corresponding bifurcation diagrams for the system (1) shown in Figures 4(a) and 4(c) for $\alpha$ varying together in the same wide range $[0,500]$ yield relatively typical bifurcation diagrams with periodic windows, period doubling bifurcations, and regions of allegedly chaotic behavior. Moreover, the diagrams for $x$ and $y$ seem to be practically identical.
@@
Figure 4. Bifurcation diagrams of variables $x$ and $y$ for the model (2) with fixed parameters $\varepsilon=0.7$, $\beta=0.2$, $n=4$ for $\alpha_{1}=\alpha_{2}$ (a), (c) and $\alpha_{2}=200$ (b), (d).
In an attempt to seek the reasons for the $x$-$y$ symmetry in the bifurcation diagrams shown in Figures 4(a) and 4(c), as well as the odd bifurcation observed in Figures 4(b) and 4(d) at $\alpha_{1}=200$, we used numerical simulations to compute the actual attractors in the phase space for several values of $\alpha_{1}$. It turns out that for $\alpha_{1}=\alpha_{2}=200$ one can see two attractors symmetric to each other, as shown in Figure 5. When $\alpha_{1}$ is increased or decreased by certain amount, while keeping the other parameters unchanged, including $\alpha_{2}=200$, one of the attractors disappears and the system exhibits a single attractor. The symmetry of the equations (1) and the existence of the two attractors explains the fact that we observed identical bifurcation diagrams in Figures 4(a) and 4(c). The disappearance of one of the attractors, on the other hand, explains the sudden change in the shape of the observed projections of the attractors in Figures 4(b) and 4(d) when $\alpha_{1}$ crosses the value of $\alpha_{2}=200$.
Figure 5. Two attractors observed in the system (2) with the parameters $\varepsilon=0.7$, $\beta=0.2$, $n=4$, and $\alpha_{1}=\alpha_{2}=200$.
@@ 7. Numerical procedure for the computation of the maximum Lyapunov exponent
Chaos is characterized by aperiodic long-term behavior of a bounded deterministic system that exhibits sensitive dependence on initial conditions. In such a system, small differences in initial conditions can lead to vastly different outcomes, making prediction difficult. A key tool for quantifying the separation rate of such trajectories are the Lyapunov exponents. In general, the existence of a positive Lyapunov exponent indicates instability and chaotic behavior, while having all the Lyapunov exponents negative suggests convergence of nearby trajectories and thus stability. In an $n$-dimensional system, there are $n$ Lyapunov exponents, corresponding to the system’s degrees of freedom. See e.g. [Spr03] for a comprehensive introduction to this topic.
Let us explain the background and the numerical approach that we use to compute the the maximum Lyapunov exponent for our system
@@
In order to compute the maximum Lyapunov exponent for the system (2), the Jacobian matrix of the right-hand side is first determined as follows:
@@
Then an initial small disturbance is iterated along the trajectory of the system using the Jacobian matrix evaluated at consecutive points of the trajectory. After each iteration, the size of the disturbance is measured. It shows how quickly nearby trajectories diverge or converge. A positive value of the maximum Lyapunov exponent suggests chaotic behavior of the system, as it indicates an average exponential divergence of nearby trajectories.
@@
Let us now describe the numerical procedure that we employed to estimate the maximum Lyapunov exponent for a considerable number of points in the parameter space. The algorithm used during these calculations relies on the fact that for a given point $x$ orbiting inside an attractor, almost every vector from the neighborhood of $x$ will converge towards the direction of fastest expansion.
@@
In order to calculate the Lyapunov exponent numerically for initial conditions $x_{0}$ (preferably in the neighborhood of an attractor), a second point $x_{0}^{\prime}$ is chosen at a small distance $d_{0}=10^{-8}$ from $x_{0}$. After one iteration of the system we obtain $d_{1}=||f(x_{0})-f(x_{0}^{\prime})||$. From the definition of the Lyapunov exponents, we expect $d_{1}\approx d_{0}e^{\lambda}$, allowing us to estimate $\lambda\approx\ln\frac{d_{1}}{d_{0}}$. This process is repeated iteratively with $x_{m+1}=f(x_{m})$ and $x_{m+1}^{\prime}=x_{m+1}+\frac{d_{0}}{d_{m+1}}(f(x_{m}^{\prime})-x_{m+1})$. Normalization is applied at each step to keep both trajectories close enough that $f$ can be reasonably approximated by its Jacobian, yet sufficiently separated to minimize numerical errors that would make their images merge into one number. The final estimate is the average of the $\lambda$ values obtained in all the iterations.
@@
Note that we treat the system as if it were effectively one-dimensional, because the vector $x_{m}^{\prime}-x_{m}$ naturally converges towards the direction of fastest expansion, which aligns with the eigenvector associated with the largest eigenvalue of the Jacobian $J$. This approach does not yield the full Lyapunov spectrum. However, it is inexpensive computationally and the estimated maximal exponent suffices for identifying chaotic dynamics in the system.
@@
If a global attractor exists, the method should converge to a single Lyapunov exponent value, independent of the chosen initial condition. This behavior has been confirmed in the current setup, as shown in Figure 6. The left panel displays the distribution of Lyapunov exponent values calculated for an array of initial conditions, showing a normal distribution centered around approximately $0.35$, which supports convergence toward a single value. The right panel shows the spatial distribution of these values over the phase space, where the uniformity and homogeneity of the values provide further evidence of its independence of the choice of $x_{0}$ and $y_{0}$. Together, these visualizations provide a baseline for understanding the computational methods used in further chapters.
@@
8. A search for positive Lyapunov exponents in both systems
The purpose of our first numerical experiment that involves the computation of the maximum Lyapunov exponent is check whether adopting different reaction rates $\alpha_{1}$ and $\alpha_{2}$ for each gene proposed in the system (2), as opposed to using the same rate $\alpha=\alpha_{1}=\alpha_{2}$ in the system (1), increases the number of parameter pairs $(\beta,\varepsilon)$ for which a positive maximum Lyapunov exponent can be encountered.
For that end, we took a lattice of $100\times 100$ different values of the parameters $\beta\in[0,1)$ and $\varepsilon\in[0,1]$. Specifically, we examined $100$ evenly spaced values of $\beta$, ranging from 00 to $0.99$, including both endpoints, and $100$ values of $\varepsilon$ ranging from 00 to $1$, inclusive. For each of these combinations, we computed the maximum Lyapunov exponent for a lattice of $1000\times 1000$ values of $\alpha_{1},\alpha_{2}\in[0,100]$, including both 0 and 100, taking the initial condition $x_{0}=0.1$, $y_{0}=0.1$ in each case. We conducted the computations for each of the two models (1) and (2) and for $n\in\{3,4\}$, which resulted in conducting four series of computations in total.
The plots in Figure 7 illustrate those pairs of parameters $\beta$ and $\varepsilon$ for which we were able to find reaction rates $(\alpha_{1},\alpha_{2})\in[0,100]\times[0,100]$ in the case of the model (2), or a reaction rate $\alpha\in[0,100]$ in the case of the model (1), with a positive maximum Lyapunov exponent. Note that the presence of the positive Lyapunov exponent indicates chaotic behavior encountered in the system.
The plots show that a larger number of parameter pairs $(\beta,\varepsilon)$ for the model (2) were found to be capable of experiencing chaotic behavior, compared to the model (1). This suggests that the model (2) is more prone to experience chaotic dynamics.

### v2

@@ 4. An absorbing set for the model
Following the ideas from [17], we shall prove that if $x_{t}\geq b_{1}$ then $x_{t+1}\leq x_{t}-c_{1}$, that is, after each iteration, the $x$ coordinate of the point decreases at least by the constant amount $c_{1}$, as long as it is beyond $b_{1}$. This statement implies that after a finite number $N_{0}$ of iterations, we must reach the condition $x_{t+N_{0}}\leq b_{1}$. Moreover, we shall also show that if $x_{t}\leq b_{1}$ then also $x_{t+1}\leq b_{1}$.
@@
(3) $\begin{split}x_{t+1}&=\alpha_{1}/(1+\text{a non-negative term})+\beta x_{t}\leq\alpha_{1}+\beta x_{t}\\ &\leq(1-\beta)x_{t}-c_{1}+\beta x_{t}=x_{t}-c_{1}.\end{split}$
Let us now consider the second statement and take $x_{t}\leq b_{1}=(\alpha_{1}+c_{1})/(1-\beta)$. Then we calculate:
(4) $\begin{split}x_{t+1}&\leq\alpha_{1}+\beta x_{t}\leq\alpha_{1}+\beta(\alpha_{1}+c_{1})/(1-\beta)\\ &=\alpha_{1}(1-\beta)/(1-\beta)+\beta(\alpha_{1}+c_{1})/(1-\beta)\\ &=(\alpha_{1}-\beta\alpha_{1}+\beta\alpha_{1}+\beta c_{1})/(1-\beta)\\ &=(\alpha_{1}+\beta c_{1})/(1-\beta)\leq(\alpha_{1}+c_{1})/(1-\beta)=b_{1}.\end{split}$
@@
In our case, it follows from Proposition 4.2 that if $\alpha_{i}\in[0,100]$ and $\beta\leq 0.5$ then all the recurrent dynamics of the model we consider is contained in $[0,200]^{2}$.
@@ 5. Periodic and chaotic behavior of the model
Both models (1) and (2) admit periodic solutions and chaotic dynamics. However, allowing $\alpha_{1}\neq\alpha_{2}$ may switch the model to quantitatively different dynamics by changing $\alpha_{2}$ alone while leaving $\alpha_{1}$ unchanged, as we illustrate below.
The plots in Figure 3 show evolution of variables $x$ and $y$ in time in model (2) for two specific trajectories, both starting at $x_{0}=y_{0}=0.5$: one for the model with $\alpha_{1}=\alpha_{2}=20$, the other one with the same value of the first reaction rate $\alpha_{1}=20$, but with $\alpha_{2}=60$. In this example, the same values of the other parameters are used in both cases: $\varepsilon=0.8$, $\beta=0.1$, $n=4$. For $\alpha_{1}=\alpha_{2}=20$ one can immediately notice the chaotic nature of the model. However, the trajectory for $\alpha_{2}=60$ appears to approach a stable periodic orbit.
The opposite situation is also present in the model. Specifically, it is possible to find a case in which for $\alpha_{1}=\alpha_{2}$ we observe periodic solutions, while changing the value of $\alpha_{2}$ yields chaotic behavior. An example of such a case occurs for the parameters $\alpha_{1}=3$, $\alpha_{2}=3$, $\beta=1$, $\varepsilon=0.8$, and $n=4$. Periodic trajectories switch to chaotic solutions when we change the value of $\alpha_{2}$ to $28$.
These plots highlight the role of $\alpha_{1}$ and $\alpha_{2}$ in controlling the model’s stability, and illustrate how the gene expression model can shift between oscillatory and chaotic regimes, depending solely on the choice of these parameters.
@@ 6. Analysis of bifurcation diagrams
A bifurcation diagram is a visual representation of the evolution of the global attractor (or attractors in general) of a dynamical system as a selected one-dimensional parameter varies. The horizontal axis of the diagram typically represents the parameter being varied, while the vertical axis corresponds to the one-dimensional phase space of the system, or a projection onto a selected variable if the actual dimension of the phase space is higher than one. Several features of the global attractor can be easily spotted on a bifurcation diagram, such as the size and complexity of the attractor, as well as bifurcations it undergoes when the parameter is varied.
The purpose of the experiment described below is to investigate the dynamics by means of bifurcation diagrams with respect to one of the reaction rates $\alpha_{1}$ while keeping the other rate $\alpha_{2}$ fixed in model (2) and compare it to the dynamics of model (1) in which both reaction rates are the same. For that purpose, we fix $\varepsilon=0.7$, $\beta=0.2$, $n=4$, $\alpha_{2}=200$, and create bifurcation diagrams for $\alpha_{1}\in[0,500]$ for model (2). We do the same for model (1) with $\alpha\in[0,500]$.
The results of this experiment are shown in Figure 4. In the vertical axes of the bifurcation diagrams, we show projections of the observed dynamics to variables $x$ and $y$, respectively. In the horizontal axes, we indicate the parameters of the models that are varied: the common reaction rate $\alpha$ in model (1) in Figures 4(a) and 4(c), and the reaction rate $\alpha_{1}$ of model (2) in Figures 4(b) and 4(d).
It is interesting to see that there is a sudden transition in model (2) that occurs in the bifurcation diagrams 4(b) and 4(d) as the parameter $\alpha_{1}$ crosses the value $\alpha_{1}=\alpha_{2}=200$. Indeed, this is kind of a boundary between two different behaviors of the model. Moreover, we notice considerable difference between the shape of diagrams 4(b) and 4(d). The first one resembles a series of period-doubling bifurcations, while the second one looks like showing period-halving bifurcations.
On the other hand, the corresponding bifurcation diagrams for model (1) shown in Figures 4(a) and 4(c) for $\alpha$ varying together in the same wide range $[0,500]$ yield relatively typical bifurcation diagrams with periodic windows, period-doubling bifurcations, and regions of allegedly chaotic behavior. Moreover, the diagrams for $x$ and $y$ seem to be practically identical.
@@
Figure 4. Bifurcation diagrams of variables $x$ and $y$ for model (2) with fixed parameters $\varepsilon=0.7$, $\beta=0.2$, $n=4$ for $\alpha_{1}=\alpha_{2}$ (a), (c) and $\alpha_{2}=200$ (b), (d).
In an attempt to seek the reasons for the $x$-$y$ symmetry in the bifurcation diagrams shown in Figures 4(a) and 4(c), as well as the odd bifurcation observed in Figures 4(b) and 4(d) at $\alpha_{1}=200$, we used numerical simulations to compute the actual attractors in the phase space for several values of $\alpha_{1}$. It turns out that for $\alpha_{1}=\alpha_{2}=200$ one can see two attractors symmetric to each other, as shown in Figure 5. When we increase or decrease $\alpha_{1}$ by a certain amount, while keeping the other parameters unchanged, including $\alpha_{2}=200$, one of the attractors disappears and the model exhibits a single attractor. The symmetry of equations (1) and the existence of the two attractors explain the fact that we observed identical bifurcation diagrams in Figures 4(a) and 4(c). The disappearance of one of the attractors, on the other hand, explains the sudden change in the shape of the observed projections of the attractors in Figures 4(b) and 4(d) when $\alpha_{1}$ crosses the value of $\alpha_{2}=200$.
Figure 5. Two attractors observed in model (2) with the parameters $\varepsilon=0.7$, $\beta=0.2$, $n=4$ and $\alpha_{1}=\alpha_{2}=200$.
@@ 7. Numerical procedure for the computation of the maximum Lyapunov exponent
Chaos is characterized by aperiodic long-term behavior of a bounded deterministic system that exhibits sensitive dependence on initial conditions. In such a system, small differences in initial conditions can lead to vastly different outcomes, making prediction difficult. A key tool for quantifying the separation rate of such trajectories is the Lyapunov exponent. In general, the existence of a positive Lyapunov exponent indicates instability and chaotic behavior, while having all the Lyapunov exponents negative suggests convergence of nearby trajectories and thus stability. In an $n$-dimensional system, there are $n$ Lyapunov exponents corresponding to the system’s degrees of freedom. See, e.g., [21] for a comprehensive introduction to this topic.
Let us explain the background and the numerical approach that we use to compute the maximum Lyapunov exponent for our model.
@@
In order to compute the maximum Lyapunov exponent for model (2), the Jacobian matrix of the right-hand side is determined as follows:
@@
Then an initial small disturbance is iterated along the trajectory, following the equations, using the Jacobian matrix evaluated at consecutive points of the trajectory. After each iteration, the size of the disturbance is measured. It shows how quickly nearby trajectories diverge or converge. A positive value of the maximum Lyapunov exponent suggests chaotic behavior of the model, as it indicates an average exponential divergence of nearby trajectories.
@@
Let us now describe the numerical procedure that we employed to estimate the maximum Lyapunov exponent for a considerable number of points in the parameter space. The algorithm used during these calculations relies on the fact that for a given point $x$ orbiting inside an attractor, almost every vector in the neighborhood of $x$ will converge towards the direction of the fastest expansion.
@@
In order to calculate the Lyapunov exponent numerically for initial conditions $x_{0}$ (preferably in a neighborhood of an attractor), a second point $x_{0}^{\prime}$ is chosen at a small distance $d_{0}=10^{-8}$ from $x_{0}$. After one iteration, we obtain $d_{1}=\|f(x_{0})-f(x_{0}^{\prime})\|$. From the definition of the Lyapunov exponents, we expect $d_{1}\approx d_{0}e^{\lambda}$, allowing us to estimate $\lambda\approx\ln\frac{d_{1}}{d_{0}}$. This process is repeated iteratively with $x_{m+1}=f(x_{m})$ and $x_{m+1}^{\prime}=x_{m+1}+\frac{d_{0}}{d_{m+1}}(f(x_{m}^{\prime})-x_{m+1})$. Normalization is applied at each step to keep both trajectories close enough so that $f$ can be reasonably approximated by its Jacobian, yet sufficiently separated to minimize numerical errors that would make their images merge into one number. The final estimate is the average of the $\lambda$ values obtained in all iterations.
@@
Note that we treat the model as if it were effectively one-dimensional, because the vector $x_{m}^{\prime}-x_{m}$ naturally converges towards the direction of fastest expansion, which aligns with the eigenvector associated with the largest eigenvalue of the Jacobian matrix $J$. This approach does not yield the full Lyapunov spectrum. However, it is inexpensive computationally and the estimated maximum exponent suffices to identify chaotic dynamics in the model.
@@
If a global attractor exists, the method should converge to a single Lyapunov exponent value, independent of the chosen initial condition. This behavior has been confirmed in the current setup, as shown in Figure 6. The left panel displays the distribution of Lyapunov exponent values calculated for a large number of initial conditions, showing a normal distribution centered around approximately $0.35$, which supports convergence towards a single value. The right panel shows the spatial distribution of these values in the phase space, where the uniformity and homogeneity of the values provide further evidence of their independence with respect to the choice of $x_{0}$ and $y_{0}$. Together, these visualizations provide a baseline for understanding the computational methods discussed in the following sections.
@@
8. A search for positive Lyapunov exponents in both models
The purpose of our first numerical experiment that involves the computation of the maximum Lyapunov exponent is to check whether taking different reaction rates $\alpha_{1}$ and $\alpha_{2}$ for each gene proposed in model (2), as opposed to using the same rate $\alpha=\alpha_{1}=\alpha_{2}$ in model (1), increases the number of parameters $(\beta,\varepsilon)$ for which a positive maximum Lyapunov exponent can be encountered.
For that purpose, we took a grid of $100\times 100$ different values of the parameters $\beta\in[0,1)$ and $\varepsilon\in[0,1]$. Specifically, we examined $100$ evenly spaced values of $\beta$, ranging from 00 to $0.99$, including both endpoints, and $100$ values of $\varepsilon$ ranging from 00 to $1$, inclusive. For each of these combinations, we computed the maximum Lyapunov exponent for a grid of $1000\times 1000$ values of $\alpha_{1},\alpha_{2}\in[0,100]$, including both 00 and $100$, taking the initial condition $x_{0}=0.1$, $y_{0}=0.1$ in each case. We conducted these computations for each of the two models (1) and (2), and for $n\in\{3,4\}$, which resulted in four series of computations in total.
The plots in Figure 7 illustrate those pairs of parameters $\beta$ and $\varepsilon$ for which we were able to find reaction rates $(\alpha_{1},\alpha_{2})\in[0,100]\times[0,100]$ in the case of model (2), or a reaction rate $\alpha\in[0,100]$ in the case of model (1), with a positive maximum Lyapunov exponent. Note that the presence of a positive Lyapunov exponent indicates chaotic behavior encountered in the model.
The plots show that a larger number of parameter pairs $(\beta,\varepsilon)$ for model (2) were found to be capable of experiencing chaotic behavior, compared to model (1). This suggests that model (2) is more prone to experience chaotic dynamics.